# Inversion-DPO: Precise and Efficient Post-Training for Diffusion Models

### v2

@@ Abstract
1 Introduction
@@
Diffusion models (DMs) [48, 46, 7, 6, 53, 76, 32, 43, 21] have achieved state-of-the-art performance in image generation. This advancement is largely attributed to recent developments in alignment methods of DMs [22, 55, 28, 8, 72, 63, 29], which enable models to be fine-tuned to better align with specific standards or human preferences. These methods typically require training two models simultaneously: a base model and a reward model. While this training paradigm incentivizes superior model performance, it also introduces significant computational overhead.
@@
Diffusion-DPO [57] reparameterizes the reward score into the loss function of the base diffusion model, eliminating the need for an auxiliary model and establishing a new paradigm for DMs to learn from human preferences. However, we identify a post-training issue in Diffusion-DPO, primarily due to its approximation of the reverse process using the forward distribution during derivation. This approximation leads to suboptimal computational accuracy and inefficient training.
@@
To overcome this limitation, we draw insights from DDIM inversion [52] to establish a precise alignment framework for diffusion model fine-tuning. Our Inversion-DPO method fundamentally enhances approximation accuracy by adopting DDIM Inversion to approximate the reverse process $p_{\theta}(x_{1:T}|x_{0})$. This proposal eliminates the distributional mismatch inherent in previous approaches, improving approximation accuracy of sampling trajectories reversely from a sample $x_{0}$ to inversion samples $x_{1:T}$.
@@
Such precision further translates to substantial efficiency gains through simplification of the final loss. By leveraging the trajectory of DDIM Inversion, our method reduces the original four KL divergence terms in Diffusion-DPO’s derivation to two matching terms between aligned model predictions and reference model prediction. This enables more than 2$\times$ faster training convergence compared to baseline implementations. The efficiency improvements are particularly crucial for large-scale diffusion models like SDXL [43], where traditional approximation methods incur high computational costs.
@@
To achieve high-quality and reliable image generation, we generalize Inversion-DPO to two specific tasks: basic text-to-image generation and compositional image generation. Extensive experiments demonstrate the effectiveness of our method in both diffusion model training and inference, as well as its superior performance compared with current SOTA methods, achieving significant improvements in key evaluation metrics (mention the specific improvements, e.g., percentage gains). LABEL:fig:_teaser presents representative generated samples. Furthermore, we introduce a novel dataset with structural annotations and complex scene preference scores, enabling structured evaluation and deeper insight into Inversion-DPO’s compositional preference learning.
@@ 2 Related Work
Alignment of Diffusion Models. As generative models are widely used, aligning outputs with user preferences has become a key focus. Researchers are integrating Reinforcement Learning from Human Feedback (RLHF) [42, 2] into diffusion models to improve controllability and accuracy. A common approach involves using pre-trained or additional reward models [22, 55, 28, 8, 72], leveraging human feedback and external signals to fine-tune generation. ImageReward [63] trains a reward model based on human feedback to improve image quality and semantic alignment. DPOK [11] and DDPO [3] combine reward signals with reinforcement learning for fine-tuning diffusion models, guiding generation through real-time feedback. ReNO [10] introduces a strategy that adjusts initial noise with reward signals, bypassing model training. IterComp [71] collects and combines model preferences, using iterative feedback to enhance generative capability.
@@
The methods above typically require explicit training of reward models. Alternatively, some approaches bypass reward models and directly fine-tune generative models via reinforcement learning. [40] For example, Diffusion-DPO [57] simplifies training by using differentiable reward signals for end-to-end fine-tuning. D3PO [65] eliminates explicit reward model training by using human feedback to guide the process. A Dense Reward View [68] enhances reward fine-tuning with dense feedback at each denoising step, improving efficiency and stability. SPO [33] adjusts denoising performance at each step with a preference model and resampler. While these DPO-based methods reduce reward model overhead, they still face post-training challenges. Recent works have explored using inversion techniques to improve DPO for diffusion models. DDIM-InPO [37] employs a reparameterization approach that treats diffusion models as single-step generators, selectively fine-tuning only the latent variables most correlated with preference data. SmPO-Diffusion [38]introduces smoothed preference distributions combined with a Renoise Inversion technique, performing multiple DDIM inversion steps followed by an additional renoise step to handle preference variability. In contrast, our Inversion-DPO directly leverages complete DDIM inversion to precisely compute the intractable posterior $p_{\theta}(x_{1:T}|x_{0})$, eliminating the approximation errors inherent in previous methods and achieving a more efficient loss formulation with only two matching terms instead of four KL divergences, improving accuracy and efficiency.
@@
Compositional Image Generation. Compositional Text-to-Image Generation is a challenge, especially when involving multiple objects and complex relationships [67, 20]. While many studies have improved generative models, the sequential text format still limits results, particularly in compositional image generation. Researchers have proposed methods to enhance control and spatial awareness. For instance, Compositional Diffusion [35] and Attend-and-Excite [5] improve generation efficiency but face limitations in complex scenes. Methods like GLIGEN [30] and Ranni [13] incorporate spatial conditioning and multimodal information to improve control, though they come with high training costs. BoxDiff [62] and RealCompo [70] optimize cross-attention and balance realism and complexity but remain dependent on bounding box accuracy. MIGC [74] and MIGC++ [73] address multi-instance composition using multimodal descriptions but focus on spatial control, failing to fully address abstract semantic relationships between objects.
@@
To overcome the limitations of text formats, some studies have focused on using scene graphs for compositional generation [23, 1, 35, 67, 60, 15]. Scene graphs (SG) consist of nodes and edges representing objects and their relationships. Compositional SG2IM methods aim to generate high-quality images with multiple objects and complex relationships by better understanding and combining scene elements [12, 58, 61].
@@
SGDiff [66] enhances scene graph-based image generation by pretraining a scene graph encoder and integrating it with Stable Diffusion. SG-Adapter [51] fine-tunes Stable Diffusion to incorporate scene graph information, improving image quality and semantic consistency. R3CD [34] introduces SG Transformers to expand diffusion models, learning abstract object interactions in larger datasets. DisCo [59] combines scene graph decomposition with VAEs and diffusion models for more diverse outputs. The LAION-SG dataset [31] further improves the model’s understanding of complex scenes.
@@
3 Preliminary
3.1 Diffusion-DPO
Reinforcement Learning from Human Feedback (RLHF) typically involves two stages: training a reward model from pairwise preferences and optimizing policies via reinforcement learning (RL). Direct Preference Optimization (DPO) [45] bypasses this complexity by directly optimizing policies using preference data under a classification objective. Formally, given preference pairs $(x_{0}^{w},x_{0}^{l})$ conditioned on prompts $c$, the Bradley-Terry model defines preference likelihood as:
$p_{\text{BT}}(x_{0}^{w}\succ x_{0}^{l}|c)=\sigma\left(r(c,x_{0}^{w})-r(c,x_{0}^{l})\right),$ (1)
@@ 3.1 Diffusion-DPO
$\max_{p_{\theta}}\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0}|c)}[r(c,x_{0})]-\beta\mathbb{D}_{\text{KL}}[p_{\theta}(x_{0}|c)\|p_{\theta_{0}}(x_{0}|c)],$ (2)
@@
$p_{\theta}^{*}(x_{0}|c)\propto p_{\theta_{0}}(x_{0}|c)\exp\left(r(c,x_{0})/\beta\right).$ (3)
@@
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(c,x_{0}^{w},x_{0}^{l})}\left[\log\sigma\left(\beta\log\frac{p_{\theta}(x_{0}^{w}|c)}{p_{\theta_{0}}(x_{0}^{w}|c)}-\beta\log\frac{p_{\theta}(x_{0}^{l}|c)}{p_{\theta_{0}}(x_{0}^{l}|c)}\right)\right].$ (4)
@@
For diffusion models, the challenge lies in defining likelihoods over generated images $x_{0}$. Diffusion-DPO [57] extends DPO by leveraging the evidence lower bound (ELBO) of diffusion processes. The KL-regularized reward maximization objective becomes:
$\max_{p_{\theta}}\mathbb{E}_{p_{\theta}(x_{1:T}|c,x_{0})}\left[R(c,x_{0:T})\right]-\beta\mathbb{D}_{\text{KL}}[p_{\theta}(x_{0:T}|c)\|p_{\theta_{0}}(x_{0:T}|c)],$ (5)
@@
$(\|\epsilon^{l}-\epsilon_{\theta}(x_{t}^{l},t)\|_{2}^{2}-\|\epsilon^{l}-\epsilon_{\theta_{0}}(x_{t}^{l},t)\|_{2}^{2}))),$ (6)
@@
3.2 DDIM Inversion
In this part, we introduce DDIM Inversion which is crucial in our proposed method. In traditional diffusion models like DDPM [17], the generative process starts with Gaussian noise $x_{T}$ and gradually denoises through a reverse Markov chain to obtain a clean image $x_{0}$. Although the training procedure maximizes a variational lower bound, the sampling process is inherently stochastic, making it difficult to precisely recover the sampling path $\{x_{t}\}$ for a given $x_{0}$.
Denoising Diffusion Implicit Models (DDIM) [52] proposes a non-Markovian inference process. Given any intermediate state $x_{t}$, the trained $\epsilon_{\theta}(x_{t},t)$ estimates the corresponding denoised sample:
$\hat{x}_{0}=f_{\theta}^{(t)}(x_{t})=\frac{x_{t}-\sqrt{1-\alpha_{t}}\,\epsilon_{\theta}(x_{t},t)}{\sqrt{\alpha_{t}}},$ (7)
@@ 3.2 DDIM Inversion
$\begin{split}p_{\theta}(x_{t-1}|x_{t})&=\mathcal{N}\left(\sqrt{\alpha_{t-1}}\hat{x}_{0}\right.\\ &\quad+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\frac{\boldsymbol{x}_{t}-\sqrt{\alpha_{t}}\hat{x}_{0}}{\sqrt{1-\alpha_{t}}},\\ &\quad\left.\sigma_{t}^{2}\boldsymbol{I}\right)\end{split}$ (8)
@@
When $t=1$, $p_{\theta}(x_{0}|x_{1})=\mathcal{N}\left(\hat{x}_{0},\sigma_{1}^{2}\boldsymbol{I}\right)$. The noise term $\sigma_{t}$ can be zero during sampling, thus converting the generative process into a deterministic implicit probabilistic model. Since the update process no longer introduces additional randomness, DDIM not only allows high-quality samples to be generated in fewer steps but also renders the generative mapping invertible, which enables recovering the sampling trajectory $\{x_{1},x_{2},\ldots,x_{T}\}$ from $x_{0}$ [52].
@@
It has been observed that when the number of sampling steps is large, predicted noises are close in adjacent steps [75]. This supports the assumption that the ordinary differential equation (ODE) process can be reversed within the limit of small steps [25], which means $\epsilon_{\theta}(x_{t},t)\approx\epsilon_{\theta}(x_{t-1},t-1)$. This key property allows us to invert the DDIM generative process [9, 24]. Starting from a real sample $x_{0}$, we can backtrack to sampling its potential denoising sequence $\{x_{1},x_{2},\ldots,x_{T}\}$. The detailed inversion iteration is
$x_{t}=\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}x_{t-1}+(\sqrt{1-\alpha_{t}}-\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}-\alpha_{t}})\epsilon_{\theta}(x_{t-1},t-1)$ (9)
@@
Such iteration finally reaches $x_{T}$. Under the assumption above, with $x_{T}$ as the initial noise, a deterministic sampling process $\epsilon_{\theta}$ is likely to recover the sequence $\{x_{T-1},\ldots,x_{1}\}$ and come back to the original $x_{0}$. More advanced estimations [14, 24] give more precise recovery. This capability to recover the latent path from $x_{0}$ provides a theoretical basis for Inversion-DPO we introduce next.
4 method

### v4

@@ Abstract.
Diffusion models, Direct Preference Optimization, Post-training
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755220isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer vision
1. Introduction
Figure 1. Visual comparison of different baselines and Inversion-DPO. Each image is annotated with PickScore at the bottom right. Inversion-DPO outperforms the baselines in both human visual perception and metric evaluation.
@@ 1. Introduction
Diffusion models (DMs) (Saharia et al., 2022; Ramesh et al., 2022; Dhariwal and Nichol, 2021; Chen et al., 2024; Tewel et al., 2024; Zhou et al., 2023; Li et al., 2024c; Podell et al., 2023; Jiang et al., 2024b) have achieved state-of-the-art performance in image generation. This advancement is largely attributed to recent developments in alignment methods of DMs (Jiang et al., 2024a; Uehara et al., 2024; Li et al., 2024a; Domingo-Enrich et al., 2025; Zhang et al., 2024a; Xu et al., 2024; Li et al., 2024d), which enable models to be fine-tuned to better align with specific standards or human preferences. These methods typically require training two models simultaneously: a base model and a reward model. While this training paradigm incentivizes superior model performance, it also introduces significant computational overhead.
@@
Diffusion-DPO (Wallace et al., 2024) reparameterizes the reward score into the loss function of the base diffusion model, eliminating the need for an auxiliary model and establishing a new paradigm for DMs to learn from human preferences. However, we identify a post-training issue in Diffusion-DPO, primarily due to its approximation of the reverse process using the forward distribution during derivation. This approximation leads to suboptimal computational accuracy and inefficient training.
@@
To overcome this limitation, we draw insights from DDIM inversion (Song et al., 2020) to establish a precise alignment framework for diffusion model fine-tuning. Our Inversion-DPO method fundamentally enhances approximation accuracy by adopting DDIM Inversion to approximate the reverse process $p_{\theta}(x_{1:T}|x_{0})$. This proposal eliminates the distributional mismatch inherent in previous approaches, improving approximation accuracy of sampling trajectories reversely from a sample $x_{0}$ to inversion samples $x_{1:T}$.
@@
Such precision further translates to substantial efficiency gains through simplification of the final loss. By leveraging the trajectory of DDIM Inversion, our method reduces the original four KL divergence terms in Diffusion-DPO’s derivation to two matching terms between aligned model predictions and reference model prediction. This enables more than 2$\times$ faster training convergence compared to baseline implementations. The efficiency improvements are particularly crucial for large-scale diffusion models like SDXL (Podell et al., 2023), where traditional approximation methods incur high computational costs.
@@
To achieve high-quality and reliable image generation, we generalize Inversion-DPO to two specific tasks: basic text-to-image generation and compositional image generation. Extensive experiments demonstrate the effectiveness of our method in both diffusion model training and inference, as well as its superior performance compared with current SOTA methods, achieving significant improvements in key evaluation metrics (mention the specific improvements, e.g., percentage gains). Figure 1 presents representative generated samples. Furthermore, we introduce a novel dataset with structural annotations and complex scene preference scores, enabling structured evaluation and deeper insight into Inversion-DPO’s compositional preference learning.
@@ 2. Related Work
Alignment of Diffusion Models. As generative models are widely used, aligning outputs with user preferences has become a key focus. Researchers are integrating Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Bai et al., 2022) into diffusion models to improve controllability and accuracy. A common approach involves using pre-trained or additional reward models (Jiang et al., 2024a; Uehara et al., 2024; Li et al., 2024a; Domingo-Enrich et al., 2025; Zhang et al., 2024a), leveraging human feedback and external signals to fine-tune generation. ImageReward (Xu et al., 2024) trains a reward model based on human feedback to improve image quality and semantic alignment. DPOK (Fan et al., 2023) and DDPO (Black et al., 2024) combine reward signals with reinforcement learning for fine-tuning diffusion models, guiding generation through real-time feedback. ReNO (Eyring et al., 2024) introduces a strategy that adjusts initial noise with reward signals, bypassing model training. IterComp (Zhang et al., 2024c) collects and combines model preferences, using iterative feedback to enhance generative capability.
@@
The methods above typically require explicit training of reward models. Alternatively, some approaches bypass reward models and directly fine-tune generative models via reinforcement learning (Majumder et al., 2024). For example, Diffusion-DPO (Wallace et al., 2024) simplifies training by using differentiable reward signals for end-to-end fine-tuning. D3PO (Yang et al., 2024b) eliminates explicit reward model training by using human feedback to guide the process. A Dense Reward View (Yang et al., 2024a) enhances reward fine-tuning with dense feedback at each denoising step, improving efficiency and stability. SPO (Liang et al., 2025) adjusts denoising performance at each step with a preference model and resampler. While these DPO-based methods reduce reward model overhead, they still face post-training challenges. Recent works have explored using inversion techniques to improve DPO for diffusion models. DDIM-InPO (Lu et al., 2025a) employs a reparameterization approach that treats diffusion models as single-step generators, selectively fine-tuning only the latent variables most correlated with preference data. SmPO-Diffusion (Lu et al., 2025b)introduces smoothed preference distributions combined with a Renoise Inversion technique, performing multiple DDIM inversion steps followed by an additional renoise step to handle preference variability. Our Inversion-DPO also leverages complete DDIM inversion. Out attempt is to precisely compute the intractable posterior $p_{\theta}(x_{1:T}|x_{0})$, eliminating the approxima- tion errors inherent in previous methods and achieving a more efficient loss formulation. Our derivation results in only two matching terms instead of four KL divergences in Diffusion-DPO, improving accuracy and efficiency.
@@
Compositional Image Generation. Compositional Text-to-Image Generation is a challenge, especially when involving multiple objects and complex relationships (Yang et al., 2024c; Huang et al., 2024). While many studies have improved generative models, the sequential text format still limits results, particularly in compositional image generation. Researchers have proposed methods to enhance control and spatial awareness. For instance, Compositional Diffusion (Liu et al., 2022) and Attend-and-Excite (Chefer et al., 2023) improve generation efficiency but face limitations in complex scenes. Methods like GLIGEN (Li et al., 2023) and Ranni (Feng et al., 2024) incorporate spatial conditioning and multimodal information to improve control, though they come with high training costs. BoxDiff (Xie et al., 2023) and RealCompo (Zhang et al., 2024b) optimize cross-attention and balance realism and complexity but remain dependent on bounding box accuracy. MIGC (Zhou et al., 2024b) and MIGC++ (Zhou et al., 2024a) address multi-instance composition using multimodal descriptions but focus on spatial control, failing to fully address abstract semantic relationships between objects.
@@
To overcome the limitations of text formats, some studies have focused on using scene graphs for compositional generation (Johnson et al., 2018; Ashual and Wolf, 2019; Liu et al., 2022; Yang et al., 2024c; Wang et al., 2023; Han et al., 2024). Scene graphs (SG) consist of nodes and edges representing objects and their relationships. Compositional SG2IM methods aim to generate high-quality images with multiple objects and complex relationships by better understanding and combining scene elements (Feng et al., 2023; Wang et al., 2024a; Wu et al., 2023).
@@
SGDiff (Yang et al., 2022) enhances scene graph-based image generation by pretraining a scene graph encoder and integrating it with Stable Diffusion. SG-Adapter (Shen et al., 2024) fine-tunes Stable Diffusion to incorporate scene graph information, improving image quality and semantic consistency. R3CD (Liu and Liu, 2024) introduces SG Transformers to expand diffusion models, learning abstract object interactions in larger datasets. DisCo (Wang et al., 2024b) combines scene graph decomposition with VAEs and diffusion models for more diverse outputs. The LAION-SG dataset (Li et al., 2024b) further improves the model’s understanding of complex scenes.
@@
3. Preliminary
3.1. Diffusion-DPO
Reinforcement Learning from Human Feedback (RLHF) typically involves two stages: training a reward model from pairwise preferences and optimizing policies via reinforcement learning (RL). Direct Preference Optimization (DPO) (Rafailov et al., 2023) bypasses this complexity by directly optimizing policies using preference data under a classification objective. Formally, given preference pairs $(x_{0}^{w},x_{0}^{l})$ conditioned on prompts $c$, the Bradley-Terry model defines preference likelihood as:
(1) $p_{\text{BT}}(x_{0}^{w}\succ x_{0}^{l}|c)=\sigma\left(r(c,x_{0}^{w})-r(c,x_{0}^{l})\right),$
@@ 3.1. Diffusion-DPO
(2) $\max_{p_{\theta}}\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0}|c)}[r(c,x_{0})]-\beta\mathbb{D}_{\text{KL}}[p_{\theta}(x_{0}|c)\|p_{\theta_{0}}(x_{0}|c)],$
@@
(3) $p_{\theta}^{*}(x_{0}|c)\propto p_{\theta_{0}}(x_{0}|c)\exp\left(r(c,x_{0})/\beta\right).$
@@
(4) $\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(c,x_{0}^{w},x_{0}^{l})}\left[\log\sigma\left(\beta\log\frac{p_{\theta}(x_{0}^{w}|c)}{p_{\theta_{0}}(x_{0}^{w}|c)}-\beta\log\frac{p_{\theta}(x_{0}^{l}|c)}{p_{\theta_{0}}(x_{0}^{l}|c)}\right)\right].$
@@
For diffusion models, the challenge lies in defining likelihoods over generated images $x_{0}$. Diffusion-DPO (Wallace et al., 2024) extends DPO by leveraging the evidence lower bound (ELBO) of diffusion processes. The KL-regularized reward maximization objective becomes:
(5) $\max_{p_{\theta}}\mathbb{E}_{p_{\theta}(x_{1:T}|c,x_{0})}\left[R(c,x_{0:T})\right]-\beta\mathbb{D}_{\text{KL}}[p_{\theta}(x_{0:T}|c)\|p_{\theta_{0}}(x_{0:T}|c)],$
@@
(6) $(\|\epsilon^{l}-\epsilon_{\theta}(x_{t}^{l},t)\|_{2}^{2}-\|\epsilon^{l}-\epsilon_{\theta_{0}}(x_{t}^{l},t)\|_{2}^{2}))),$
@@
3.2. DDIM Inversion
In this part, we introduce DDIM Inversion which is crucial in our proposed method. In traditional diffusion models like DDPM (Ho et al., 2020), the generative process starts with Gaussian noise $x_{T}$ and gradually denoises through a reverse Markov chain to obtain a clean image $x_{0}$. Although the training procedure maximizes a variational lower bound, the sampling process is inherently stochastic, making it difficult to precisely recover the sampling path $\{x_{t}\}$ for a given $x_{0}$.
Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020) proposes a non-Markovian inference process. Given any intermediate state $x_{t}$, the trained $\epsilon_{\theta}(x_{t},t)$ estimates the corresponding denoised sample:
(7) $\hat{x}_{0}=f_{\theta}^{(t)}(x_{t})=\frac{x_{t}-\sqrt{1-\alpha_{t}}\,\epsilon_{\theta}(x_{t},t)}{\sqrt{\alpha_{t}}},$
@@ 3.2. DDIM Inversion
(8) $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}\left(\sqrt{\alpha_{t-1}}\hat{x}_{0}+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\frac{\boldsymbol{x}_{t}-\sqrt{\alpha_{t}}\hat{x}_{0}}{\sqrt{1-\alpha_{t}}},\sigma_{t}^{2}\boldsymbol{I}\right)$
@@
When $t=1$, $p_{\theta}(x_{0}|x_{1})=\mathcal{N}\left(\hat{x}_{0},\sigma_{1}^{2}\boldsymbol{I}\right)$. The noise term $\sigma_{t}$ can be zero during sampling, thus converting the generative process into a deterministic implicit probabilistic model. Since the update process no longer introduces additional randomness, DDIM not only allows high-quality samples to be generated in fewer steps but also renders the generative mapping invertible, which enables recovering the sampling trajectory $\{x_{1},x_{2},\ldots,x_{T}\}$ from $x_{0}$ (Song et al., 2020).
@@
It has been observed that when the number of sampling steps is large, predicted noises are close in adjacent steps (Zhou et al., 2024c). This supports the assumption that the ordinary differential equation (ODE) process can be reversed within the limit of small steps (Kim et al., 2022), which means $\epsilon_{\theta}(x_{t},t)\approx\epsilon_{\theta}(x_{t-1},t-1)$. This key property allows us to invert the DDIM generative process (Dong et al., 2023; Ju et al., 2023). Starting from a real sample $x_{0}$, we can backtrack to sampling its potential denoising sequence $\{x_{1},x_{2},\ldots,x_{T}\}$. The detailed inversion iteration is
(9) $x_{t}=\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}x_{t-1}+(\sqrt{1-\alpha_{t}}-\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}-\alpha_{t}})\epsilon_{\theta}(x_{t-1},t-1)$
@@
Such iteration finally reaches $x_{T}$. Under the assumption above, with $x_{T}$ as the initial noise, a deterministic sampling process $\epsilon_{\theta}$ is likely to recover the sequence $\{x_{T-1},\ldots,x_{1}\}$ and come back to the original $x_{0}$. More advanced estimations (Garibi et al., 2025; Ju et al., 2023) give more precise recovery. This capability to recover the latent path from $x_{0}$ provides a theoretical basis for Inversion-DPO we introduce next.
4. Method